# Skein modules and character varieties of Seifert manifolds

### v1

@@ 5.3. Reducedness
It is known that each point of $X(M)$ is represented by a completely reducible representation $\rho$ [LM85, Sik12]. In particular, $\rho$ can be taken to be either irreducible or diagonal. Furthermore, the tangent space $T_{\rho}\,\mathcal{X}(M)$ at an irreducible $\rho$ is isomorphic to $H^{1}(M,\mathrm{Ad}\,\rho)$, see [LM85]. Similarly, $H^{1}(M,\mathrm{Ad}\,\rho)=0,$ for a diagonal representation $\rho$ implies that $[\rho]$ is reduced in $\mathcal{X}(M),$ by [Sik12, Theorem 1] and [HP23, Lemma 21]. Therefore, Theorem 5.9 follows from Lemmas 5.11, 5.12, 5.13 and 5.14 below.
@@
The proof is similar to the argument in [BC06, Lemma 2.4]: Since $\rho$ is irreducible, $B^{1}(M,\mathrm{\rho})$ is of dimension $3,$ so we only need to show that $Z^{1}(M,\mathrm{Ad}\,\rho)=3$. We use again the presentation of $\pi_{1}(M)$ given in Equation 8. Since $h,c_{1},c_{2},c_{3}$ generate $\pi_{1}(M)$, the cocycle condition implies that any $\varepsilon\in Z^{1}(M,\mathrm{Ad}\,\rho)$ is determined by $H:=\varepsilon(h)$ and $X_{i}:=\varepsilon(c_{i}),$ for $i=1,2,3$. Furthermore, by (16) we obtain $\varepsilon(c_{i}^{-1})=-\mathrm{Ad}\,\rho(c_{i}^{-1})\varepsilon(c_{i})$. Hence applying $\varepsilon$ to the relations $[c_{i},h]=1$, and utilizing properties (13), (15), (16), we obtain
@@
Since $\rho(c_{i})$ is of finite order for each $i$, it is diagonalizable with eigenvalues $\zeta_{i},\zeta^{-1}_{i}$ such that $\zeta_{i}^{p_{i}}=(\pm 1)^{q_{i}}$. We note that $\zeta_{i}\neq\pm 1$ for each $i$. Indeed, if say $\zeta_{1}=\pm 1$, then the image of $\rho$ is contained in the set of powers of ±$\pm\rho(c_{2})$ implying that $\rho$ is abelian. In a basis of a diagonalization of $\rho(c_{i}),$ $\mathrm{Ad}\,\rho(c_{i})$ sends $X_{i}=\begin{pmatrix}x_{i}&y_{i}\\ z_{i}&-x_{i}\end{pmatrix}\in sl_{2}(\mathbb{C})$ to $\begin{pmatrix}x_{i}&\zeta_{i}^{2}y_{i}\\ \zeta_{i}^{-2}z_{i}&-x_{i}\end{pmatrix}$, for $i=1,2,3$. Since $\zeta_{i}^{2p_{i}}=1$, we get $01+\zeta_{i}^{2}+\ldots+\zeta_{i}^{2p_{i}-2}=$ 0, which implies that that $\left(I+\mathrm{Ad}\,\rho(c_{i})+\ldots+\mathrm{Ad}\,\rho(c_{i}^{p_{i}-1})\right)X_{i}$ is diagonal. Therefore, Equation (20) is satisfied if and only if $X_{i}$ has diagonal zero, which is equivalent to $\operatorname{Tr}(\rho(c_{i})X_{i})=0$.
@@
Applying $\varepsilon$ to the equations $[c_{i},h]=1$ we obtain (17) which now reduces to
@@
(24) $-\left(\frac{q_{1}}{p_{1}}+\frac{q_{2}}{p_{2}}+\frac{q_{3}}{p_{3}}\right)u=-e(M)u=0.$
@@
If $\rho$ maps two of $c_{1},c_{2},c_{3}$ to ±$\pm I$ then so it does the third one and $\rho$ is central. Hence, $\rho$ is non-exceptional and the statement follows from Lemma 5.11. Therefore, we assume without loss of generality, that either $\rho(c_{1})=\pm I$ and $\rho(c_{2})\neq\pm I$, or that $\rho$ is exceptional.
Similarly to the proof of Lemma 5.13, we have $\dim B^{1}(M,\mathrm{Ad}\,\rho)=2$ and we compute $\dim Z^{1}(M,\mathrm{Ad}\,\rho)$. As before, $\rho(c_{i})=\mathrm{Diag}(\zeta_{i},\zeta_{i}^{-1})$, for $i=1,2,3$. We also keep the notations for the entries of $H,X_{1},X_{2},X_{3}$. As in the proof of Lemma 5.12, applying $\varepsilon$ to relation $[c_{2},h]=1$ gives $\mathrm{Ad}\,\rho(c_{2})(H)-H=0$. Hence $H$ commutes with $\rho(c_{2})$ and thus it is diagonal, i.e. $v=w=0$. (Now the $1$-cocycle condition is satisfied for $[c_{i},h]=1$, $i=1$ and $3$, as well.)
@@
If $\rho(c_{i})\neq\pm I$ then (19) is automatically satisfied on off-diagonal entries. However, if $\rho(c_{1})=\pm I$ then (19), and the diagonality of $H$ implies $y_{1}=z_{1}=0$. Hence, the $1$-cocycle conditions for the relations $[c_{i},h]=1$ and $c_{i}^{p_{i}}h^{q_{i}}=1$, define a $7$-dimensional space of cocycles (determined by parameters: $u,y_{i},z_{i}$ for $i=1,2,3$) for exceptional $\rho$’s and a $5$-dimensional space of cocycles (determined by $u,y_{i},z_{i}$ for $i=2,3$) for non-exceptional $\rho$’s.
@@ 5.4. Bases for $\mathbb{C}[\mathcal{X}(M)]$ and $\mathcal{S}(M,\mathbb{Q}(A))$
Let $S_{-1}(M):=S(M,\mathbb{Z}[A^{\pm 1}])\otimes_{\mathbb{Z}[A^{\pm 1}]}\mathbb{C},$ where the $\mathbb{Z}[A^{\pm 1}]$-module structure of $\mathbb{C}$ is given by sending $A$ to $-1$. By Przytycki-Sikora [PS00], $S_{-1}(M)$ has the structure of a $\mathbb{C}$-algebra that is isomorphic to the coordinate ring $\mathbb{C}[\mathcal{X}(M)]$ of $\mathcal{X}(M)$, through the isomorphism $\psi:S_{-1}(M)\to\mathbb{C}[\mathcal{X}(M)]$, sending any framed link $L=L_{1}\cup\ldots\cup L_{k}$ in $M$ to $(-1)^{k}t_{L}$ where $t_{L}=t_{L_{1}}\cdot\ldots\cdot t_{L_{k}}$. For another approach, up to nilpotents, see [Bul97]. Here, $t_{L_{i}}$ is the trace function of $L_{i}$ with its framing ignored. As we explained in [DKS23, Proposition 3.3], if $S(M,\mathbb{Q}[A^{\pm 1}])$ is tame and $\mathcal{X}(M)$ is reduced, a basis of $S_{-1}(M)\simeq\mathbb{C}[\mathcal{X}(M)]$ leads to a basis of $\mathcal{S}(M)=\mathcal{S}(M,\mathbb{Q}(A))$.
@@
Let $\mathbb{K}$ be a field of characteristic zero and let $P\in\mathbb{K}[X_{1},\ldots,X_{n}]$ be a polynomial of degree $\leq d_{i}$ in the variable $X_{i},$ for $i=1,...,n$. If $P$ vanishes on a subset of $\mathbb{K}^{n}$ of the form $S_{1}\times\ldots\times S_{n},$ where $|S_{i}|=d_{i}+1,$ then $P=0.$
@@
Proof of Theorem 5.15. The last claim of the theorem follows from the first part by [DKS23, Proposition 3.3(b)]. Let us prove the first part.
@@
since the other components of $F$ vanish on $X_{2,\tau}(M)$. Now, recall from the proof of Proposition 5.8 that $X_{2}(M)$ contains characters taking all possible values $(\chi(c_{1}),\chi(c_{2}),\chi(c_{3})\in C_{p_{1}}^{+}\times C_{p_{2}}^{+}\times C_{p_{3}}^{+}$. In other words, characters in $X_{+,\tau}(M)$ take all possible values $(\chi(c_{2}),\chi(c_{3}))$ in $C_{p_{2}}^{+}\times C_{p_{3}}^{+}$ for every $\tau\in C_{p_{1}}^{+}.$ Since for each $\tau\in C_{p_{1}}^{+},$ (25) is a polynomial of degree $<p_{i}^{+}=|C_{p_{i}}^{+}|$, for $i=2,3,$ which vanishes on $C_{p_{2}}^{+}\times C_{p_{3}}^{+}$,
@@
Since $F$ vanishes for $y_{M}=2$ as well, assume $y_{M}>2$ now. Then $G:=\frac{F}{t_{h}^{2}-4}$ is a polynomial in $t_{h}$ of degree $y_{M}-3+\delta_{M}$ that vanishes on all of the abelian characters of $H_{1}(M)$ for which we have $\chi(h)\neq\pm 2$. By Lemma 5.17, $H_{1}(M)$ is generated by $h$ and, hence, $t_{h}$ takes $y_{M}$ distinct values on abelian characters, including $y_{M}-2+\delta_{M}$ values which are not ±$\pm 2$. This implies that all coefficients of $G$ vanish, i.e. $\nu_{j}=0$ for every $j.$

### v2

@@ 5.3. Reducedness
It is known that each point of $X(M)$ is represented by a completely reducible representation $\rho$ [LM85, Sik12]. In particular, $\rho$ can be taken to be either irreducible or diagonal. Furthermore, the tangent space $T_{\rho}\,\mathcal{X}(M)$ at an irreducible $\rho$ is isomorphic to $H^{1}(M,\mathrm{Ad}\,\rho)$, see [LM85]. Similarly, $H^{1}(M,\mathrm{Ad}\,\rho)=0,$ for a diagonal representation $\rho$ implies that $\rho]$ is reduced in $\mathcal{X}(M),$ by [Sik12, Theorem 1] and [HP23, Lemma 21]. Therefore, Theorem 5.9 follows from Lemmas 5.11, 5.12, 5.13 and 5.14 below.
@@
The proof is similar to the argument in [BC06, Lemma 2.4]: Since $\rho$ is irreducible, $B^{1}(M,\mathrm{\rho})$ is of dimension $3,$ so we only need to show that $Z^{1}(M,\mathrm{Ad}\,\rho)=3$. We use again the presentation of $\pi_{1}(M)$ given in Equation 8. Since $h,c_{1},c_{2},c_{3}$ generate $\pi_{1}(M)$, the cocycle condition implies that any $\varepsilon\in Z^{1}(M,\mathrm{Ad}\,\rho)$ is determined by $H:=\varepsilon(h)$ and $X_{i}:=\varepsilon(c_{i}),$ for $i=1,2,3$. Furthermore, by (16) we obtain $\varepsilon(c_{i}^{-1})=-\mathrm{Ad}\,\rho(c_{i}^{-1})\varepsilon(c_{i})$. Hence applying $\varepsilon$ to the relations $c_{i},h]=1$, and utilizing properties (13), (15), (16), we obtain
@@
Since $\rho(c_{i})$ is of finite order for each $i$, it is diagonalizable with eigenvalues $\zeta_{i},\zeta^{-1}_{i}$ such that $\zeta_{i}^{p_{i}}=(\pm 1)^{q_{i}}$. We note that $\zeta_{i}\neq\pm 1$ for each $i$. Indeed, if say $\zeta_{1}=\pm 1$, then the image of $\rho$ is contained in the set of powers of $\pm\rho(c_{2})$ implying that $\rho$ is abelian. In a basis of a diagonalization of $\rho(c_{i}),$ $\mathrm{Ad}\,\rho(c_{i})$ sends $X_{i}=\begin{pmatrix}x_{i}&y_{i}\\ z_{i}&-x_{i}\end{pmatrix}\in sl_{2}(\mathbb{C})$ to $\begin{pmatrix}x_{i}&\zeta_{i}^{2}y_{i}\\ \zeta_{i}^{-2}z_{i}&-x_{i}\end{pmatrix}$, for $i=1,2,3$. Since $\zeta_{i}^{2p_{i}}=1$, we get $1+\zeta_{i}^{2}+\ldots+\zeta_{i}^{2p_{i}-2}=0$, which implies that that $\left(I+\mathrm{Ad}\,\rho(c_{i})+\ldots+\mathrm{Ad}\,\rho(c_{i}^{p_{i}-1})\right)X_{i}$ is diagonal. Therefore, Equation (20) is satisfied if and only if $X_{i}$ has diagonal zero, which is equivalent to $\operatorname{Tr}(\rho(c_{i})X_{i})=0$.
@@
Applying $\varepsilon$ to the equations $c_{i},h]=1$ we obtain (17) which now reduces to
@@
(24) $\left(\frac{q_{1}}{p_{1}}+\frac{q_{2}}{p_{2}}+\frac{q_{3}}{p_{3}}\right)u=-e(M)u=0.$
@@
If $\rho$ maps two of $c_{1},c_{2},c_{3}$ to $\pm I$ then so it does the third one and $\rho$ is central. Hence, $\rho$ is non-exceptional and the statement follows from Lemma 5.11. Therefore, we assume without loss of generality, that either $\rho(c_{1})=\pm I$ and $\rho(c_{2})\neq\pm I$, or that $\rho$ is exceptional.
Similarly to the proof of Lemma 5.13, we have $\dim B^{1}(M,\mathrm{Ad}\,\rho)=2$ and we compute $\dim Z^{1}(M,\mathrm{Ad}\,\rho)$. As before, $\rho(c_{i})=\mathrm{Diag}(\zeta_{i},\zeta_{i}^{-1})$, for $i=1,2,3$. We also keep the notations for the entries of $H,X_{1},X_{2},X_{3}$. As in the proof of Lemma 5.12, applying $\varepsilon$ to relation $c_{2},h]=1$ gives $\mathrm{Ad}\,\rho(c_{2})(H)-H=0$. Hence $H$ commutes with $\rho(c_{2})$ and thus it is diagonal, i.e. $v=w=0$. (Now the $1$-cocycle condition is satisfied for $c_{i},h]=1$, $i=1$ and $3$, as well.)
@@
If $\rho(c_{i})\neq\pm I$ then (19) is automatically satisfied on off-diagonal entries. However, if $\rho(c_{1})=\pm I$ then (19), and the diagonality of $H$ implies $y_{1}=z_{1}=0$. Hence, the $1$-cocycle conditions for the relations $c_{i},h]=1$ and $c_{i}^{p_{i}}h^{q_{i}}=1$, define a $7$-dimensional space of cocycles (determined by parameters: $u,y_{i},z_{i}$ for $i=1,2,3$) for exceptional $\rho$’s and a $5$-dimensional space of cocycles (determined by $u,y_{i},z_{i}$ for $i=2,3$) for non-exceptional $\rho$’s.
@@ 5.4. Bases for $\mathbb{C}[\mathcal{X}(M)]$ and $\mathcal{S}(M,\mathbb{Q}(A))$
Let $S_{-1}(M):=S(M,\mathbb{Z}[A^{\pm 1}])\otimes_{\mathbb{Z}[A^{\pm 1}]}\mathbb{C},$ where the $\mathbb{Z}[A^{\pm 1}]$-module structure of $\mathbb{C}$ is given by sending $A$ to 1-1. By Przytycki-Sikora [PS00], $S_{-1}(M)$ has the structure of a $\mathbb{C}$-algebra that is isomorphic to the coordinate ring $\mathbb{C}[\mathcal{X}(M)]$ of $\mathcal{X}(M)$, through the isomorphism $\psi:S_{-1}(M)\to\mathbb{C}[\mathcal{X}(M)]$, sending any framed link $L=L_{1}\cup\ldots\cup L_{k}$ in $M$ to $(-1)^{k}t_{L}$ where $t_{L}=t_{L_{1}}\cdot\ldots\cdot t_{L_{k}}$. For another approach, up to nilpotents, see [Bul97]. Here, $t_{L_{i}}$ is the trace function of $L_{i}$ with its framing ignored. As we explained in [DKS25, Proposition 3.3], if $S(M,\mathbb{Q}[A^{\pm 1}])$ is tame and $\mathcal{X}(M)$ is reduced, a basis of $S_{-1}(M)\simeq\mathbb{C}[\mathcal{X}(M)]$ leads to a basis of $\mathcal{S}(M)=\mathcal{S}(M,\mathbb{Q}(A))$.
@@
Let $\mathbb{K}$ be a field of characteristic zero and let $P\in\mathbb{K}[X_{1},\ldots,X_{n}]$ be a polynomial of degree $\leq d_{i}$ in the variable $X_{i},$ for i=1,,ni=1,...,n. If $P$ vanishes on a subset of $\mathbb{K}^{n}$ of the form $S_{1}\times\ldots\times S_{n},$ where $|S_{i}|=d_{i}+1,$ then $P=0.$
@@
Proof of Theorem 5.15. The last claim of the theorem follows from the first part by [DKS25, Proposition 3.3(b)]. Let us prove the first part.
@@
since the other components of $F$ vanish on $X_{2,\tau}(M)$. Now, recall from the proof of Proposition 5.8 that $X_{2}(M)$ contains characters taking all possible values $(\chi(c_{1}),\chi(c_{2}),\chi(c_{3})\in C_{p_{1}}^{+}\times C_{p_{2}}^{+}\times C_{p_{3}}^{+}$. In other words, characters in $X_{+,\tau}(M)$ take all possible values $(\chi(c_{2}),\chi(c_{3}))$ in $C_{p_{2}}^{+}\times C_{p_{3}}^{+}$ for every $\tau\in C_{p_{1}}^{+}.$ Since for each $\tau\in C_{p_{1}}^{+},$ (25) is a polynomial of degree $p_{i}^{+}=|C_{p_{i}}^{+}|$, for $i=2,3,$ which vanishes on $C_{p_{2}}^{+}\times C_{p_{3}}^{+}$,
@@
Since $F$ vanishes for $y_{M}=2$ as well, assume $y_{M}>2$ now. Then $G:=\frac{F}{t_{h}^{2}-4}$ is a polynomial in $t_{h}$ of degree $y_{M}-3+\delta_{M}$ that vanishes on all of the abelian characters of $H_{1}(M)$ for which we have $\chi(h)\neq\pm 2$. By Lemma 5.17, $H_{1}(M)$ is generated by $h$ and, hence, $t_{h}$ takes $y_{M}$ distinct values on abelian characters, including $y_{M}-2+\delta_{M}$ values which are not $\pm 2$. This implies that all coefficients of $G$ vanish, i.e. $\nu_{j}=0$ for every $j.$